Two machines make identical parts which are combined on a production
line. The older machine makes 40% of the units, of which 85% are of
satisfactory quality. The newer machine makes 60% of the units, of which
92% are satisfactory. A random check further down the production line
shows an unusual fault which suggests the machine which made the unit
needs some adjustment. Which machine is more likely to have made the
unit?

Answer by stanbon(75887)

Two machines make identical parts which are combined on a production
line. The older machine makes 40% of the units, of which 85% are of
satisfactory quality.
Older machine DATA:
unsatisfactory final product = 0.14*0.4 = 0.056 or 5.6%
-------------------------------------

The newer machine makes 60% of the units, of which
92% are satisfactory.
Newer machine DATA:
unsatisfactory final product = 0.08*0.6 = 0.048 or 4.8%
-------------------------------------
A random check further down the production line
shows an unusual fault which suggests the machine which made the unit
needs some adjustment. Which machine is more likely to have made the
unit?
Ans: Looks like the older machine is more likely to be the culprit.
-----
Cheers,
Stan H.